# Manipulation Priors Learned From Egocentric Videos

### v1

@@ 3.2 Encoding Manipulation Priors
We aim to train a visual encoder tailored for dexterous manipulation by reasoning about human hand-object interactions extracted from egocentric videos. We use an encoder-decoder structure as a stepping stone in learning useful representations: the encoder processes input prediction frames into embedding vectors, which the decoder then uses to predict contact points and future hand poses. The decoder relies entirely on the encoder’s embeddings, which have a much lower dimensionality than the input images, to infer its output. This bottleneck design encourages the encoder to produce information tailored specifically for object manipulation-related predictions, rather than learning to produce “general-purpose” visual embeddings [45, 4].
Contact Point Prediction. We employ the cross-entropy loss to learn to predict contact points $pt_{p}$ in the prediction frame $f_{p}$. To do so, we discretize the image’s width $w$ and height $h$ into $B_{x}$ resp.  $B_{y}$ bins each and let the decoder predict the respective bin indices into which the contact point should fall. We set $B_{x}=100$ and $B_{y}=100$ in our experiments. We formulate the contact loss $\mathcal{L}_{con,pt}$ for one sample and one contact point $pt$ as the following:
$\mathcal{L}_{con,pt_{p}}=CE({\arg\max}_{j\in\{1,...,B_{x}\}}\ \tilde{p}_{pt,x,j},\left\lfloor\frac{\hat{x}}{w}\times B_{x}\right\rfloor)+$
@@
where $\tilde{p}_{pt,x}\in\mathbb{R}^{B_{x}}$, $\tilde{p}_{pt,y}\in\mathbb{R}^{B_{y}}$ are the decoder’s corresponding logit vectors. $CE(\tilde{c},\hat{c})$ represents the cross-entropy loss for the predicted class $\tilde{c}$ and the ground-truth class $\hat{c}$. $\hat{x}$ and $\hat{y}$ are the extracted contact point coordinates. We use $pt\in\{0,1\}$ for the thumb and index contact points.
Hand Pose Prediction. During MAPLE’s training, we let the decoder predict a probability distribution for each hand pose token in the token sequence, then use the highest-scoring token logit to determine the final token from which to reconstruct the hand pose. Similarly to the contact point prediction, we also use a cross-entropy loss for this. Specifically, we process a MANO hand pose $\mathcal{H}\in\mathbb{R}^{21\times 3}$ using a tokenization producedure
@@
$\mathcal{L}_{hand}=\sum_{n=1}^{N}CE({\arg\max}_{j\in\{1,...,C\}}\ \tilde{p}_{n,j},t_{n}).$
Ensuring Temporal Distinctiveness. We hypothesize that the temporal distinctiveness of the input features plays an important role in the success of policy learning. Note that our initial weights are copied from DINO [4], which provides a highly distinctive visual prior due to DINO’s contrastive training objectives. To maintain this distinctiveness, we can optimize over a small set of parameters, e.g. those of the LayerNorm [25] modules of the ViT [13] backbone, which would change the representation less aggressively and preserve the prior distinctiveness. Alternatively, we can employ a temporal contrastive loss [41] during training to maintain distinctiveness in the face of more drastic changes. Both approaches have been explored in prior work [55, 40] and are incorporated in MAPLE as MAPLE-LN, where we train only the LayerNorm parameters, and MAPLE-AP, where all parameters are trained.
For a given frame $f^{v_{m}}_{j}$ at position $j$ of video $v_{m}$, we follow the InfoNCE objective employed in [40], pushing representations of frames from different videos ($f^{v_{m}}_{j}$ and $f^{v_{n}}_{\ell}$, $v_{m}\neq v_{n}$) or farther in time ($f^{v_{m}}_{i}$ and $f^{v_{m}}_{k}$, $i<j<k$) apart while pulling representations from close frames in the same video ($f^{v_{m}}_{i}$ and $f^{v_{m}}_{j}$, $i<j$) together:
$\mathcal{L}_{ctr}=-\log\Bigg{(}\frac{e^{s(f^{v_{m}}_{i},f^{v_{m}}_{j})}}{e^{s(f^{v_{m}}_{i},f^{v_{m}}_{j})}+e^{s(f^{v_{m}}_{i},f^{v_{m}}_{k})}+e^{s(f^{v_{m}}_{j},f^{v_{n}}_{\ell})}}\Bigg{)}.$
$s$ is a function increasing with the similarity of its input vectors, such as $s(x,y)=-\lVert x-y\rVert^{2}_{2}$.
@@
$\mathcal{L}=\sum_{pt\in\{0,1\}}\mathcal{L}_{con,pt_{p}}+\lambda_{hand}\mathcal{L}_{hand}+\lambda_{ctr}\mathcal{L}_{ctr}.$
We simply set $\lambda_{ctr}=1$ for MAPLE-AP and $\lambda_{ctr}=0$ for MAPLE-LN, as well as $\lambda_{hand}=1$ for both models.
Implementation Details. Similar to other ViT [13]-derived baselines [55, 32, 4], we employ a ViT-B/16 as the encoder’s architecture. The encoder is initialized with the weights of [4] and produces visual features of dimension $d=768$. To ensure fair comparisons and evaluate the contribution of our supervised training procedure, rather than the choice of architecture, we do not opt for a larger or more intricate backbone. We use a Transformer [62] decoder model with 2 Transformer layers and 4 attention heads per layer. The encoder and decoder are jointly trained using the AdamW [30] optimizer with a learning rate of $5\times 10^{-5}$. We use a batch size of 128 when training MAPLE-LN, and 512 when training MAPLE-AP.
@@ 4 Experiments
In our experiments, we evaluate the effectiveness of MAPLE for robotic dexterous manipulation tasks in both simulation and real-world environments.
@@
Figure 4: Simulated Evaluation Environments. We evaluate our method on four environments from DAPG [47] (a-d) and propose three new robotic environments (e-g). Our new environments aim to evaluate the manipulation capabilities of a set of objects commonly used by humans, namely a brush, a clothes iron, and a pan.
4.1 Experimental Setup
Training Data. We run our extraction pipeline on Ego4D [15], which consists of more than 3,000 hours of unscripted human activities across diverse scenarios (e.g., gardening, exercise, and cooking). To ensure a fair comparison, we follow the established protocols as in [40, 55] and select the same subset covering approximately 1,590 hours of video for extracting data. Our pipeline produces approximately 82,100 training samples from this subset. See the Supp. Mat. for more details.
@@ 4.1 Experimental Setup
Baseline Encoders. We evaluate against both the commonly used, general-purpose visual encoder DINO [4], as well as four state-of-the-art (SotA) encoders designed specifically for robotic manipulation:
@@
  • DINO [4]: A self-supervised visual encoder trained with a contrastive loss on images and their extracted patches.
@@
  • R3M [40]: A robotic manipulation baseline using a temporal contrastive loss on images and another contrastive loss estimating task completion based on video summaries matched with visual features for pairs of images.
@@
  • HRP [55]: A visual encoder trained for robotic manipulation settings using contact points extracted from hand-object bounding box intersections and wrist trajectories extracted using FrankMocap [51].
@@
  • VC-1 [32]: A visual encoder targeting a variety of embodied intelligence tasks, trained on large amounts of images using the Masked Auto-Encoder [20] paradigm.
  • VIP [31]: A visual encoder for robotic manipulation using a novel implicit time contrastive objective.
@@
For all methods based on ViT [13], namely DINO, HRP, and MAPLE, we employ a ViT-B/16 [13] backbone. R3M and VIP use a ResNet50 [19] backbone. We initialize all baselines with their publicly available weights.
@@ 4.2 Evaluation in Simulation Environments
Existing Simulation Environments and Tasks. We first evaluate on the four tasks from the established DAPG [47] benchmark suite. The tasks consist of open door, use hammer, rotate pen, and relocate ball (see Figure 4).
New Environment and Task Designs. We further introduce a new set of four custom-designed, challenging evaluation tasks focused on dexterous manipulation that require more fine-grained control. Specifically, we design three tasks in the MuJoCo simulator [59] that require complex dexterous manipulation of tools using an Adroit hand [24]. All three tasks feature objects standing on a table, with the Adroit hand hovering above them and able to move freely (see Figure 4 for visualization). These tasks are (1) Pan: Lift a pan and place it on the induction plate of the nearby stove; (2) Brush: Retrieve the brush from a cup, then stroke it onto the nearby canvas; and (3) Clothes Iron: Grasp a clothes iron on a table by the handle and slide it along the leg of the pants positioned nearby. Detailed examples and renderings of the tasks are in Supp. Mat. Sec. S2.2.
Collect Demonstrations for the New Tasks. To conform with the established evaluation standard in literature [40, 31], we record 25 expert demonstrations for each of our tasks using Rokoko Smartgloves [49], employing the algorithm from [44] to retarget the human hand pose to the Adroit hand in the simulation. Please refer to the Supp. Mat. for details regarding the demonstration collection and the goal definitions of each task.
Model Door Hammer Pen Relocate Brush Iron Pan Mean
DINO [4] 19.6 $\pm$ 5.1 20.0 $\pm$ 0.9 60.4 $\pm$ 1.7 28.9 $\pm$ 0.8 5.6 $\pm$ 2.2 8.0 $\pm$ 1.9 4.2 $\pm$ 1.7 21.0
HRP [55] 17.3 $\pm$ 3.9 16.2 $\pm$ 2.5 61.8 $\pm$ 4.1 29.8 $\pm$ 0.8 4.2 $\pm$ 0.3 1.6 $\pm$ 1.1 1.1 $\pm$ 1.6 18.9
VC-1 [32] 14.7 $\pm$ 1.9 24.4 $\pm$ 4.6 56.7 $\pm$ 5.2 23.6 $\pm$ 0.8 4.4 $\pm$ 0.8 3.3 $\pm$ 0.5 0.4 $\pm$ 0.3 18.2
R3M [40] 15.3 $\pm$ 2.0 12.9 $\pm$ 3.6 60.0 $\pm$ 0.5 40.4 $\pm$ 1.3 7.3 $\pm$ 2.0 15.1 $\pm$ 1.1 0.2 $\pm$ 0.3 21.6
VIP [31] 11.1 $\pm$ 2.7 18.6 $\pm$ 4.3 66.4 $\pm$ 1.7 19.8 $\pm$ 1.3 5.1 $\pm$ 0.8 4.2 $\pm$ 1.4 2.4 $\pm$ 0.3 18.2
@@
MAPLE-AP 22.0 $\pm$ 3.5 18.2 $\pm$ 4.9 63.8 $\pm$ 0.8 29.6 $\pm$ 3.0 4.2 $\pm$ 1.3 5.8 $\pm$ 2.3 0.9 $\pm$ 0.6 20.6
Table 1: Results on Dexterous Simulation Environments. We report the task completion success rate for seven simulation tasks, as well as the mean success rate. MAPLE-LN reaches the best mean performance on the evaluation suite, underscoring its excellent generalization ability. Green indicates the best performance, and blue the second best. We report the standard deviations across randomization seeds.
Experiment Protocol. Our evaluation uses both existing tasks as well as our newly designed ones. To evaluate a given visual encoder on a given task, MLP-based policies [40] are trained to control a robot hand in simulation given proprioceptive input and features from the visual encoder. Policies are evaluated every 1,000 out of a total of 20,000 training steps, with each evaluation consisting of 50 rollouts in the simulator. During each rollout, we randomize the position and/or rotation of the object(s). The policy receives a binary score indicating whether it fulfilled the intended task according to an environment-specific success criterion, e.g. opening a door. The score of the encoder on the task is then the best average success rate across all 20 evaluations. Following prior work R3M [40], we calculate scores of encoders on tasks by additionally averaging the scores across three camera views, and three random seeds per view used to vary the policy network’s initialization, for a total of nine experiments per encoder and task. This averaging procedure helps improve the statistical robustness of the evaluation. For our custom-designed tasks, we employ a temporal horizon of 750 time steps. Gaussian noise is additionally added to the generated action to simulate imperfect execution and increase the tasks’ difficulty. See the Supp. Mat. for success definitions on our tasks.
Results. Table 1 compares our method MAPLE with various baselines and SotA methods on seven different dexterous manipulation tasks that require various levels of fine-grained control. We observe that MAPLE exhibits a strong generalization ability across tasks, and achieves the best mean performance score of all methods. MAPLE-LN outperforms MAPLE-AP for most tasks. The mean performance of R3M [40] is close to ours. However, it heavily relies on in-house human-annotated language captions during training (with a strong benefit in Relocate), while our method can learn in an entirely self-supervised manner. Surprisingly, the contrastive training objective of DINO [4] appears to give a strong performance, probably due to its strong semantic features that are robust to occlusion. The general low performance observed on our custom-designed tasks highlights the inherent complexity and challenges associated with tasks that require more complex dexterous skills and validates the usefulness of our newly introduced evaluation tasks.
$\mathcal{L}_{con,pt}$ $\mathcal{L}_{hand}$ DAPG MetaWorld FK
44.0 73.6 55.5
44.3 78.2 52.5
$\bar{\mathrm{T}}$ 44.0 81.6 55.2
$\Box$ 44.4 77.3 50.7
$1$ 46.3 77.6 51.6
@@
Table 2: Ablation of Loss Terms on Dexterous and Non-Dexterous Environments. We report the results of training our MAPLE-LN model. ✓indicates the inclusion, while ✗indicates the exclusion of a given loss term. The $\mathcal{L}_{hand}$ loss without the hand tokenizer is indicated by $\bar{\mathrm{T}}$. We further ablate sampling contact points randomly from hand-object box intersections ($\Box$) or using only one contact prediction point (1).
Ablation Studies. We ablate the contribution of our contact and hand pose losses on the tasks used by [40] in Table 2, including DAPG (pen and relocate), as well as from the non-dexterous MetaWorld (5 tasks) [17, 67], and Franka Kitchen (5 tasks) [17] benchmarks.
Dexterous Tasks Benefit From $\mathcal{L}_{hand}$ And Its Tokenization. The dexterous tasks as well as the MetaWorld tasks benefit significantly from the inclusion of our hand pose loss, as evidenced by comparing the second row with the last row. Compared to a setting where the decoder reconstructs raw hand joint configurations, using the hand pose tokenizer is helpful for the dexterous setting while harming the performance in the non-dexterous settings.
The Benefits of $\mathcal{L}_{con,pt}$. We notice that randomly sampling points from the intersection of hand and object bounding boxes hurts performance compared to our usage of the projected fingertips. Furthermore, reasoning about a single contact point (index finger only) for $\mathcal{L}_{con,pt}$ gives a worse performance on two out of three benchmark suites than reasoning about the index and thumb finger jointly.

### v2

@@ 3.3 Encoding Manipulation Priors
We aim to train a visual encoder tailored for dexterous manipulation by reasoning about human hand-object interactions extracted from egocentric videos. We use an encoder-decoder structure as a stepping stone in learning useful representations: the encoder processes input prediction frames into embedding vectors, which the decoder then uses to predict contact points and future hand poses. The decoder relies entirely on the encoder’s embeddings, which have a much lower dimensionality than the input images, to infer its output. This bottleneck design encourages the encoder to produce information tailored specifically for object manipulation-related predictions, rather than learning to produce “general-purpose” visual embeddings [CLIP, DINO].
Contact Point Prediction. We employ the cross-entropy loss to learn to predict contact points $pt_{p}$ in the prediction frame $f_{p}$. To do so, we discretize the image’s width $w$ and height $h$ into $B_{x}$ resp.  $B_{y}$ bins each and let the decoder predict the respective bin indices into which the contact point should fall. We set $B_{x}=100$ and $B_{y}=100$ in our experiments. We formulate the contact loss $\mathcal{L}_{ct,pt}$ for one sample and one contact point $pt$ as the following:
$\mathcal{L}_{ct,pt}=CE({\arg\max}_{j\in\{1,...,B_{x}\}}\ \tilde{p}_{pt,x,j},\left\lfloor\frac{\hat{x}}{w}\times B_{x}\right\rfloor)+$
@@
where $\tilde{p}_{x,\hat{x}}\in\mathbb{R}^{B_{x}}$, $\tilde{p}_{y,\hat{y}}\in\mathbb{R}^{B_{y}}$ are the decoder’s corresponding logit vectors. CE(c~,c^)CE($\tilde$$\hat{c}$) represents the cross-entropy loss for the predicted class $\tilde{c}$ and the ground-truth class $\hat{c}$. $\hat{x}$ and $\hat{y}$ are the extracted contact point coordinates. We use $pt\in\{0,1\}$ for the thumb and index contact points.
Hand Pose Prediction. During MAPLE’s training, we let the decoder predict a probability distribution for each hand pose token in the token sequence and use the highest-scoring token logit to determine the final token from which to reconstruct the hand pose. Similarly to the contact point prediction, we also use a cross-entropy loss for this. Specifically, we process a MANO hand pose $\mathcal{H}\in\mathbb{R}^{21\times 3}$ using a tokenization producedure
@@
$\mathcal{L}_{hand}=\sum_{in=1}^{N}CE({\arg\max}_{j\in\{1,...,C\}}\ \tilde{p}_{n,j},p_{n}).$
@@
$\mathcal{L}=\sum_{pt\in\{0,1\}}\mathcal{L}_{ct,pt}+\lambda_{hand}\mathcal{L}_{hand}$
3.4 Training Strategy
Training Data. We run our extraction pipeline on Ego4D [Ego4D], which consists of more than 3,000 hours of unscripted human activities across diverse scenarios (e.g., gardening, exercise, and cooking). To ensure a fair comparison, we follow the established protocols as in [R3M, HRP] and select the same subset covering approximately 1,590 hours of video for extracting data. Our pipeline produces approximately 82,100 training samples from this subset. See the Supp. Mat. for more details and visualizations.
Implementation Details. Similar to other ViT [Dosovitskiy20ViT]-derived baselines [HRP, VC1, DINO], we employ a ViT-B/16 as the encoder’s architecture. The encoder is initialized with the weights of [DINO] and produces visual features of dimension $d=768$. To ensure fair comparisons and evaluate the performance of our supervised training procedure, rather than the choice of architecture, we do not opt for a larger or more intricate backbone. We use a Transformer [Vaswani17Transformer] decoder model with 2 Transformer layers and 4 attention heads per layer. The encoder and decoder are jointly trained using the AdamW [AdamW] optimizer with a learning rate of $5\times 10^{-5}$. We use a batch size of 128 when training MAPLE.
@@ 4 Experiments
To demonstrate the generality of MAPLE, we evaluate on both perception-oriented contact prediction tasks (Sec. 4.1) and robotic dexterous manipulation tasks, which are tested in simulation (Sec. 4.2) and a real-world setup (Sec. 4.3).
@@
Figure 4: Simulated Evaluation Environments. We evaluate our method on four environments from DAPG [DAPG] (a-d) and propose four new robotic environments (e-h). Our new environments aim to evaluate the manipulation capabilities of a set of objects commonly used by humans, namely a pan, a brush, a drill, and a clothes iron.
11footnotetext:  We compare with open-source models and do not include models that do not have open access.
@@
Baseline Encoders. We evaluate against both the commonly used, general-purpose visual encoders CLIP [CLIP] and DINOv1/v2/v3 [DINO, DINOv2, DINOv3], as well as three state-of-the-art (SotA) encoders designed specifically for robotic manipulation. The baselines can be described as follows 1:
  • CLIP [CLIP]: A visual encoder trained jointly with a language encoder in a representation alignment task on a large number of text-image pairs from the internet.
@@
  • DINO [DINO]: A self-supervised visual encoder trained with a contrastive loss on images and their extracted patches. We evaluate on the three released versions DINO [DINO], DINOv2 [DINOv2] and DINOv3 [DINOv3].
@@
  • R3M [R3M]: A robotic manipulation baseline using a temporal contrastive loss on images and another contrastive loss estimating task completion based on video summaries matched with visual features for pairs of images.
@@
  • HRP [HRP]: A visual encoder trained for robotic manipulation settings using contact points extracted from hand-object bounding box intersections and wrist trajectories extracted using FrankMocap [FrankMocap].
@@
  • VC-1 [VC1]: A visual encoder targeting a variety of embodied intelligence tasks, trained on large amounts of images using the Masked Auto-Encoder [MAE] paradigm.
@@
For the methods based on ViT [Dosovitskiy20ViT], namely CLIP, DINO, DINOv3, HRP, VC-1 and MAPLE, we employ a ViT-B/16 [Dosovitskiy20ViT] backbone, with DINOv2 using ViT-B/14 by its authors’ design. R3M uses a ResNet50 [ResNet] backbone. We initialize all baselines with their publicly available weights.
Result Summary. We evaluate MAPLE on (1) contact anticipation, (2) simulated dexterous manipulation across 8 tasks including 4 new benchmarks, (3) real-world dexterous tasks using an ORCA [ORCA] hand. Our results show that MAPLE consistently improves generalization and fine-grained control. Specifically, MAPLE achieves more accurate contact localization, leading to improved first contact and stronger implicit grasp priors, which yield greater success of in-hand manipulation. Its prior-infused features consistently outperform generic representations.
4.1 Evaluation of Contact Point Prediction
We first compare MAPLE with other baseline encoders on the contact prediction task. Similar to the setup in [liu2022joint], we predict future contact points given a single input frame, whereas the original task formulation takes video sequences as input. This task has applications in robotics [RoboABC, DenseMatcher, Anygrasp] and object interaction anticipation [ContextAwareAnticipation, TransFusion, Ego4D]. We train cVAE-based contact prediction heads on features extracted from all baselines and our encoder. Following the original protocol, we evaluate Affordance Map Similarity (SIM) and Normalized Scanpath Saliency (NSS) on the dataset introduced in [liu2022joint]. This dataset is derived from various egocentric cooking datasets [EK55, EK100, EGTEAGazePlus] and provides 34.5K interaction hotspots, which we use as contact prediction ground-truth points. As evident from Table 1, MAPLE outperforms all baselines in both metrics, confirming the usefulness of its learned features. Note that these frame-based predictions are different from the results reported in the original paper, which shows video-based predictions. See Supp. Mat. for metric definitions and qualitative examples.
Table 1: Contact Prediction Evaluation. Comparison of various vision encoders on the dataset introduced in [liu2022joint]. MAPLE achieves the best performance on both SIM and NSS metrics.
Metric CLIP DINO DINOv2 DINOv3 HRP VC-1 R3M MAPLE
SIM $\uparrow$ 0.051 0.061 0.061 0.059 0.061 0.064 0.068 0.068
NSS $\uparrow$ 0.561 0.577 0.647 0.615 0.602 0.637 0.651 0.654
@@ 4.2 Evaluation in Simulation Environments
Existing Simulation Environments and Tasks. We first evaluate on the four dexterous tasks from the established DAPG [DAPG] benchmark suite, including open door, use hammer, rotate pen, and relocate ball (see Figure 4).
New Environment and Task Designs. We further introduce a new set of four custom-designed, challenging evaluation tasks focused on dexterous manipulation that require more fine-grained control. Specifically, we design four tasks in the MuJoCo simulator [MuJoCo] that require complex dexterous manipulation of tools using an Adroit hand [AdroitHand]. All four tasks feature objects standing on a table, with the Adroit hand hovering above them and able to move freely (see Figure 4 for visualization). These tasks are (1) Pan: Lift a pan and place it on the induction plate of the nearby stove; (2) Brush: Retrieve the brush from a cup, then stroke it onto the nearby canvas; (3) Drill: Lift an electric drill and align the drill bit to strike a wooden board while keeping the finger on the drill’s trigger switch; and (4) Iron: Grasp a clothes iron on a table by the handle and slide it along the leg of the pants positioned nearby. Further details and visualizations of the tasks are in Supp. Mat.
Collecting Demonstrations for the New Tasks. To conform with the established evaluation standard in literature [R3M], we record 25 expert demonstrations for each of our tasks using Rokoko Smartgloves [Smartglove], employing the algorithm from [DexMV] to retarget the human hand to the Adroit hand in the simulation. Please refer to the Supp. Mat. for the details and goal definitions of each task.
Table 2: Results on Simulation Environments. We report the task completion success rate for eight simulation tasks. Green indicates the best performance, and blue the second best. We report the standard deviations across randomization seeds. MAPLE reaches the best mean performance on the evaluation suite with variance reduced across tasks with higher complexity.
Model Door Hammer Pen Relocate Brush Drill Iron Pan Mean
@@
[…]
Experiment Protocol. Our evaluation uses both existing tasks and newly designed ones. To evaluate a given visual encoder on a given task, MLP-based policies [R3M] are trained to control a robot hand in simulation given proprioceptive input and features from the visual encoder. Policies are evaluated every 1,000 out of a total of 20,000 training steps, with each evaluation consisting of 50 rollouts in the simulator. During each rollout, we randomize the position and/or rotation of the object(s). The policy receives a binary score indicating whether it fulfilled the intended task according to an environment-specific success criterion, e.g. opening the door by a wide enough angle. The score of the encoder on the task is then the best average success rate across all 20 evaluations. Following prior work R3M [R3M], we calculate scores of encoders on tasks by additionally averaging the scores across three camera views, and three random seeds per view used to vary the policy network’s initialization, for a total of nine experiments per encoder and task. This averaging procedure helps improve the statistical robustness of the evaluation. For our custom-designed tasks, we employ a temporal horizon of 750 time steps. Gaussian noise is additionally added to the generated action to simulate imperfect execution and increase the tasks’ difficulty. See the Supp. Mat. for success definitions.
Results. Table 2 compares MAPLE with various baselines and SotA methods across eight dexterous manipulation tasks. MAPLE achieves the best overall performance among all methods, exceeding the second-best approach by more than two standard deviations. While other baselines occasionally outperform our method on individual tasks, MAPLE exhibits strong generalization across all 8 tasks, achieving the highest success rate in four tasks and ranking a close second-best in a fifth, whereas no other considered method leads in more than one.
Except for MAPLE, encoders explicitly designed for robotic manipulation struggle to outperform the generalist DINO and CLIP models in terms of mean success rates, indicating a lack of generalizability beyond their training domains. In contrast, MAPLE excels in fine-grained manipulation tasks such as the in-hand pen reorientation, brush and drill tasks, all of which require precise grasping.
Finally, the high success rates of MAPLE in our newly introduced environments, despite heavily randomized initial tool positions, further validate its robustness and efficacy for object localization and manipulation.
Table 3: Ablation of Loss Terms. We report the effect of training with and without the contact loss, as well as different types of hand pose losses on the Door, Hammer, Pen and Relocate tasks. Checkmarks (✓) indicate the inclusion and crosses (✗) indicate the exclusion of a loss term. The use of the hand pose loss without the hand tokenizer is indicated by $\bar{\mathrm{T}}$. Success rates are averages across all seeds and camera views of the respective tasks.
$\mathcal{L}_{ct}$ $\mathcal{L}_{hand}$ Door Hammer Pen Relocate Mean
29.3 10.0 72.0 32.7 36.0
32.4 9.8 68.9 31.1 35.6
$\bar{\mathrm{T}}$ 31.8 13.3 70.2 31.6 36.7
31.8 13.3 75.1 32.9 38.3
Analysis. We ablate the contribution of our contact and hand pose losses on the four DAPG dexterous tasks [DAPG] in Table 3. Performance on the in-place, yet highly dexterous pen task suffers from excluding $\mathcal{L}_{hand}$ while being less sensitive to the exclusion of the localization-focused $\mathcal{L}_{ct}$, as expected from the task’s in-place nature. In all settings, training the model with a hand pose tokenizer leads to a higher performance than letting it reconstruct hand poses from scratch. The less dexterous but localization-focused door task, requiring the policy to locate the handle and push it down in any feasible way, suffers most from the exclusion of $\mathcal{L}_{ct}$. This shows the benefits of using hand pose tokenization and demonstrates the usefulness of $\mathcal{L}_{hand}$ for dexterous tasks as well as $\mathcal{L}_{ct}$ for tasks emphasizing localization.